# How and Why LLMs Use Deprecated APIs in Code Completion? An Empirical Study

### v2

@@ III-C LLM-based Code Completion
We leverage multiple LLMs in the code completion task to observe their performance. The LLMs include both open-source and closed-source models, whose parameter sizes ranges from 350 million to 175 billion. The complete list of the LLMs are presented in Table II.
  • CodeGen-350m, 2b, 6b: CodeGen [3] is a family of large language models developed by Salesforce specifically for code generation. These models are trained on diverse programming languages and are designed to assist in writing code by providing intelligent code suggestions and completions.
  • DeepSeek-1.3b: DeepSeek-Coder [6] is designed on top of advanced transformer-based models tailored for code-related applications. This model is trained on a diverse set of coding repositories, which include a variety of programming languages and coding styles. It combines state-of-the-art machine learning techniques to provide robust code suggestions and completions.
@@
  • StarCoder2-3b: StarCoder2 [4] is an advanced language model optimized for coding tasks. It leverages large-scale pre-training on code datasets to understand programming languages deeply. StarCoder2 excels in code completion, bug detection, and code translation, making it a valuable tool for developers seeking to enhance their coding efficiency and accuracy.
  • CodeLlama-7b: CodeLlama [5] is a specialized variant of Meta’s LLaMA architecture [54], adapted for programming tasks. It is designed to support developers by providing code suggestions, completing code snippets, and assisting in debugging. CodeLlama is trained on a vast corpus of programming languages and can perform a variety of code-related tasks with high proficiency.
@@
  • GPT-3.5: GPT-3.5 [55] is a general-purpose language model developed by OpenAI. While it is not exclusively designed for coding, it possesses powerful code generation capabilities due to its extensive training on diverse text, including programming languages. GPT-3.5 can understand and generate code across various languages, making it a versatile tool for code completion, documentation, and coding assistance. When used for code-related tasks, an instruction (prompt) is often provided to guide the model to generate the desired code output.
@@
In our implementation, the first six code LLMs were downloaded from Hugging Face [56]. For LLMs with Python-specific versions available, we utilized those versions to improve completion results for Python functions. Consequently, the versions used for CodeGen and CodeLlama were CodeGen-{350M,2B,6B}-mono and CodeLlama-7b-Python-hf, which were fine-tuned on additional Python corpora. For DeepSeek and StarCoder2, the versions employed were deepseek-coder-1.3b-instruct and starcoder2-3b, respectively. For the general purpose LLM, i.e., GPT-3.5, we queried the model via its official online APIs [57], using the gpt-3.5-turbo version released in January 2024.
@@
For the six LLMs specifically tailored for code-related tasks, i.e., three versions of CodeGen, DeepSeek, StarCoder2, and CodeLlama, the constructed prompts can be directly fed into the models to generate completions. Meanwhile, for the general-purpose GPT-3.5, we provided an instruction to ensure accurate code completion for the constructed prompts. The instruction was: “Complete and output the next line for the following Python function: $pmpt$. For all these LLMs, we utilized greedy decoding (i.e., choosing the token with highest possibility at each decoding step) to generate one completion for each prompt in $\mathcal{O}$ or $\mathcal{U}$. The maximal output token limit was set to 50. The greedy decoding for GPT-3.5 was implemented through setting the temperature parameter to 0.
Formally, the procedure of LLM-based completion is defined as:
@@
TABLE II: Evaluated Large Language Models (LLMs).
  Model # Parameters Open-Source
  CodeGen-350m 350 million
CodeGen-2b 2 billion
CodeGen-6b 6 billion
  DeepSeek-1.3b 1.3 billion
  StarCoder2-3b 3 billion
  CodeLlama-7b 7 billion
  GPT-3.5 175 billion
@@ III-D Completion Result Annotation
For each sample ($pmpt$, $dep\to rep$), we examined the completions generated by LLMs and determine whether the studied API was predicted and whether the deprecated API or replacing API was predicted. Specifically, we extract the FQN of the API invocation in the predicted line using the same object type resolution and API alias resolution in Sec. III-B1. The annotation procedure is formally described as follows, $comp$ is the completion result:
$\{\emph{good},\emph{bad},\emph{irrelevant}\}\leftarrow\text{annotate}(comp)$
Specifically, bad denotes that the LLM gives a completion suggestion using a deprecated API, i.e., the FQN of an invocating API was matched to $dep$; good denotes the LLM gives a completion suggestion using a replacing API, i.e., the FQN of an invocating API was matched to $rep$; irrelevant denotes that the LLM suggests neither of the mapping APIs. Moreover, if a completion was annotated as either bad or good, we treat it as plausible. This indicates that the LLM successfully understood the code context and selected a plausible API functionality.
@@
  • API Usage Plausibility (AUP): This metric measures the portion of plausible completions, which were annotated as good or bad. AUP is defined as:
    $AUP=\frac{1}{|\mathcal{P}|}\sum_{pmpt\in\mathcal{P}}\mathbb{I}(\text{annotate}(\textbf{LLM}(pmpt))\in\{\emph{good},\emph{bad}\})$
  • Deprecated Usage Rate (DUR): This metric calculates the rate of plausible completions that were annotated as bad. DUR is defined as:
    $DUR=\frac{\sum_{pmpt\in\mathcal{P}}\mathbb{I}(\text{annotate}(\textbf{LLM}(pmpt))=\emph{bad})}{\sum_{pmpt\in\mathcal{P}}\mathbb{I}(\text{annotate}(\textbf{LLM}(pmpt))\in\{\emph{good},\emph{bad}\})}$
In these equations, $\mathcal{P}$ is the prompt set corresponding to $\mathcal{O}$ or $\mathcal{U}$, and $\mathbb{I}(\cdot)$ is a binary function that returns 1 if the passed argument is true and 0 otherwise.
@@ IV Study Results
IV-A RQ1: Model Perspective for Status Quo and Cause Analysis
@@ IV-A1 Status Quo Analysis
Figure 3 shows the distribution of good and bad completions by the different LLMs, and Table III presents the AUP and DUR metrics.
Figure 3: Distribution of good, bad, and irrelevant Completions by LLMs for Prompts from $\mathcal{O}$ and $\mathcal{U}$
TABLE III: AUP and DUR Metrics for $\mathcal{O}$, $\mathcal{U}$, and Overall Dataset (i.e., All = $\mathcal{O}\cup\mathcal{U}$)
@@
  CodeGen-350m 8.6 14.2 12.3 77.7 9.0 24.9
CodeGen-2b 19.0 25.2 23.1 89.6 11.3 32.6
CodeGen-6b 26.1 30.9 29.3 90.0 11.1 34.2
DeepSeek-1.3b 10.2 13.4 12.3 69.7 11.6 27.2
StarCoder-3b 8.8 10.9 10.2 79.4 17.0 34.4
CodeLlama-7b 27.6 32.1 30.6 86.9 12.1 34.2
GPT-3.5 13.5 16.0 15.2 85.7 17.8 37.4
API Usage Plausibility. The completion distribution and the low AUP highlight that all the LLMs faced challenges in predicting plausible API usages for the given prompts. The AUP of the LLMs for overall dataset (i.e., All $=\mathcal{O}\cup\mathcal{U}$) ranges from 10% to 30%, indicating that a majority of predictions were irrelevant. Among the LLMs, CodeLlama-7b achieved the highest AUP (30.6%), while StarCoder2-3b had the lowest overall AUP (10.2%). This may be attributed to the fact that StarCoder2 was not specifically fine-tuned on additional Python corpora, unlike other LLMs such as CodeLlama and CodeGen. Comparing the three versions of CodeGen suggests that the capacity of LLMs to predict plausible API usages increased with model size (i.e., 12.3%, 23.1%, and 29.3% for CodeGen-350m, -2b, and -6b, respectively), given the model architecture and training data remain consistent. However, it’s noteworthy that the largest LLM, GPT-3.5, did not achieve a high AUP (15.2%), possibly due to the instruction used not being finely tuned with advanced prompt engineering techniques such as chain-of-thought (COT) [58] and in-context learning (ICL) [59].
Finding 1: All the evaluated LLMs faced challenges in predicting plausible API usages, with AUP ranging from 10% to 30%. Effectiveness of the LLMs in code completion generally improved with model size and language-specific fine-tuning.
@@
Finding 2: All the evaluated LLMs faced issues with deprecated API usages, with DUR ranging from 25% to 38%, and larger models exhibiting higher DUR. Among the LLMs, CodeLlama-7b and CodeGen-6b demonstrated the best balance between AUP and DUR.
IV-A2 Root Cause Analysis
From the model perspective, the causes of deprecated API usages can be divided into two main points:
Model Training: The LLMs were trained on large-scale code corpora, primarily collected from code repositories. As libraries evolved, both deprecated APIs and their replacements were used in software development, leading to training corpora containing instances of deprecated API usages. For instance, in this study, we collected 9,022 functions from open-source code corpora that used deprecated APIs. When trained on such data, LLMs might “memorize” these deprecated APIs and their usage contexts as part of the learned knowledge [60, 61]. The different training corpora also led to the different AUP and DUR of the LLMs.
Model Inference: During the inference stage, LLMs generated completions by predicting token probabilities based on their learned prior knowledge (e.g., the memorized API usage contexts) and applying token selection strategies (e.g., greedy search or beam search). Given certain contexts, LLMs were likely to predict deprecated API usages due to the high token probabilities, without considering any posterior knowledge about API deprecations.
Finding 3: There are two primary reasons why LLMs predict deprecated APIs: the presence of deprecated API usages in corpora during model training, and the absence of posterior knowledge about API deprecations during model inference.
IV-B RQ2: Prompt Perspective for Status Quo and Cause Analysis
@@ IV-B1 Status Quo Analysis
Table III presents the AUP and DUR metrics for prompts from the two datasets, i.e., $\mathcal{O}$ and $\mathcal{U}$.
@@
API Usage Plausibility. Between the prompts from the two different datasets, $\mathcal{O}$ and $\mathcal{U}$, there are some differences in AUP for all LLMs, with relative differences (i.e., $(AUP^{\mathcal{U}}-AUP^{\mathcal{O}})/AUP^{\mathcal{O}}$) ranging from 16% to 65%. This disparity may be attributed to the imbalance in the number of outdated and up-to-dated functions in the LLMs’ training corpora [62, 63]. Indirect evidence for this is that, in this study, the up-to-dated functions collected from open-source code repositories were about twice as many as the outdated functions (i.e., 19,103 vs. 9,022), even though there was no collection preference. Given that LLMs were often trained on open-source code repositories, they likely learned more up-to-dated functions than outdated functions, leading to better AUP for the $\mathcal{U}$ dataset. Additionally, larger LLMs showcased smaller AUP differences (e.g., 16% for CodeLlama-7b), possibly due to the better generalizability.
@@
Finding 4: The LLMs showcased difference in AUP between the two datasets $\mathcal{O}$ and $\mathcal{U}$. This difference is possibly attributed to the different distribution of outdated and up-to-dated functions in the training corpora of LLMs.
@@
Deprecated Usage Rate. When considering $\mathcal{O}$ and $\mathcal{U}$ separately, all LLMs consistently demonstrated extremely high deprecated usage rates for $\mathcal{O}$ (70%-90% DUR) and relatively low rates for $\mathcal{U}$ (9%-18% DUR). This significant difference is also evident in the distribution of good and bad completions shown in Figure 3. In fact, for plausible completions, the rate of reference API usages (i.e., the usages used in the original functions) equals AUP for $\mathcal{O}$, while it is $(1-AUP)$ for $\mathcal{U}$. Considering this rate, there is no noticeable difference between $\mathcal{O}$ and $\mathcal{U}$ (i.e., around 70%-90% for both). This suggests that LLMs predicted the reference APIs for most prompts from both datasets, influenced by their differing characteristics. Nonetheless, the 9%-18% DUR for $\mathcal{U}$ indicates that LLMs still predicted the usage of deprecated APIs, even for the prompts from up-to-dated functions.
@@
Finding 5: The LLMs consistently exhibited a significant difference in DUR between the two datasets, $\mathcal{O}$ and $\mathcal{U}$, with extremely high deprecated API usage rates for $\mathcal{O}$ (70%-90% DUR) and relatively low rates for $\mathcal{U}$ (9%-18% DUR).
IV-B2 Root Cause Analysis

### v3

@@ III-C LLM-based Code Completion
We leverage multiple LLMs including open-source and closed-soure with varying parameter sizes and observe their performance on the code completion task. The complete LLM list is presented in Table II.
  • CodeGen-350m, 2b, 6b: CodeGen [3] is a family of LLMs developed by Salesforce specifically for code generation.
  • DeepSeek-1.3b: DeepSeek-Coder [6] is designed on top of Transformer and tailored for code-related applications.
@@
  • StarCoder2-3b: StarCoder2 [4] is an LLM optimized for coding tasks. It leverages large-scale pre-training on code datasets to understand programming languages deeply.
  • CodeLlama-7b: CodeLlama [5] is a specialized variant of Meta’s LLaMA [61], adapted for programming tasks.
@@
  • GPT-3.5: GPT-3.5 [62] is a general-purpose language model developed by OpenAI. While it is not exclusively designed for coding, it possesses powerful code generation capabilities due to its extensive training on diverse text, including programming languages.
@@
The first six code LLMs were downloaded from Hugging Face [63]. For LLMs with Python-specific versions available, we utilized those versions to improve completion results for Python functions. Consequently, the versions used for CodeGen and CodeLlama were CodeGen-{350M,2B,6B}-mono and CodeLlama-7b-Python-hf, which were fine-tuned on additional Python corpora. For DeepSeek and StarCoder2, the versions employed were deepseek-coder-1.3b-instructstarcoder2-3b, respectively. For the general purpose LLM, i.e., GPT-3.5, we queried the model via its official online APIs [64], using the gpt-3.5-turbo version released in January 2024.
@@
For the six LLMs specifically tailored for code-related tasks, i.e., three versions of CodeGen, DeepSeek, StarCoder2, and CodeLlama, the constructed prompts can be directly fed into the models to generate completions. Meanwhile, for the general-purpose GPT-3.5, we provided an instruction to specify the task, preventing the model from performing beyond code completion. The instruction was: “Complete and output the next line for the following Python function: $pmpt$. For all these LLMs, we utilized greedy decoding (i.e., choosing the token with highest possibility at each decoding step) to generate one completion for each prompt in $\mathcal{O}$ or $\mathcal{U}$. The maximal output token limit was set to 50. The greedy decoding for GPT-3.5 was implemented through setting the temperature parameter to 0.
The procedure of LLM-based completion is defined as:
@@
TABLE II: Evaluated Large Language Models (LLMs)
  Model # Params Python Fine-Tuned Open-Source
  CodeGen-350m 350 M
CodeGen-2b 2 B
CodeGen-6b 6 B
DeepSeek-1.3b 1.3 B
StarCoder2-3b 3 B
CodeLlama-7b 7 B
GPT-3.5 175 B UNKNOWN
@@ III-D Completion Result Annotation
For each sample ($pmpt$, $dep\to rep$), we examined the completions generated by LLMs and determine whether the studied API was predicted and whether the deprecated API or replacing API was predicted. Specifically, we extract the FQN of the API invocation in the predicted line using the same object type resolution and alias resolution in Sec. III-B1. The annotation procedure is formally described as follows:
$\{\emph{DepC},\emph{RepC},\emph{Others}\}\leftarrow\text{anno}(comp)$
Specifically, DepC denotes that the LLM gives a Deprecated Completion suggestion, and RepC denotes that the LLM gives a Replacing Completion suggestion. We identify DepC and RepC by matching the FQN of an invocating API to either $dep$ or $rep$. Others denotes that the LLM suggests neither of the mapping APIs. Moreover, if a completion was annotated as either DepC or RepC, we treat it as plausible. This indicates that the LLM successfully understood the code context and selected a plausible API functionality.
III-E Metrics
@@
  • API Usage Plausibility (AUP): This metric measures the portion of plausible completions, which were annotated as DepC or RepC. AUP is defined as:
    $AUP=\frac{1}{|\mathcal{P}|}\sum_{p\in\mathcal{P}}\mathbb{I}(\text{anno}(\textbf{LLM}(p))\in\{\emph{DepC},\emph{RepC}\})$
  • Deprecated Usage Rate (DUR): This metric calculates the rate of plausible completions that were annotated as DepC. DUR is defined as:
    $DUR=\frac{\sum_{p\in\mathcal{P}}\mathbb{I}(\text{anno}(\textbf{LLM}(p))=\emph{DepC})}{\sum_{p\in\mathcal{P}}\mathbb{I}(\text{anno}(\textbf{LLM}(p))\in\{\emph{DepC},\emph{RepC}\})}$
$\mathcal{P}$ is the prompt set (i.e., $\mathcal{O}$ or $\mathcal{U}$), and $\mathbb{I}(\cdot)$ is a binary function that returns 1 if the passed argument is true and 0 otherwise.
The API Usage Plausibility (AUP) measures how effectively LLMs predict “accurate” APIs without considering their deprecation status. This metric is crucial because merely counting deprecated API usages (DUR) does not fully capture how LLMs manage API deprecations. For instance, a low number of deprecated API usages might be due to the prevalence of Others completions, suggesting influences beyond deprecation. To provide fair comparisons across models, prompts, and libraries, we calculate the Deprecated Usage Rate (DUR) based on AUP. A balanced relationship between AUP and DUR, characterized by a relatively high AUP and low DUR, indicates that LLMs effectively predict both up-to-date and “accurate” APIs. In contrast, an imbalance suggests either a lower rate of “accurate” predictions or a prevalence of outdated APIs. All the LLMs exhibited low AUPs due to the generation of many Others predictions, likely because many APIs (e.g., PyTorch APIs) share similar usage contexts and offer flexible combinations. In addition, although these Others may provide alternative solutions for similar functionalities, they fall outside our focus on LLMs’ handling of API deprecations. To clarify, we formally defined both Others and AUP to avoid confusion with terms like “wrong” or “incorrect”. Nevertheless, we recognize the value of exploring this issue and leave a broader question for future work: how to effectively evaluate the functional correctness of generated completions, given the numerous alternative implementations for the same goal.
@@ IV Study Results
IV-A RQ1: Model Perspective for Status Quo and Root Causes
@@ IV-A1 Status Quo Analysis
Figure 3 shows the distribution of RepC and DepC completions by the different LLMs, and Table III presents the AUP and DUR metrics.
API Usage Plausibility. The completion distribution and the low AUP highlight that all the LLMs faced challenges in predicting plausible API usages for the given prompts. The AUP of the LLMs for overall dataset (i.e., All $=\mathcal{O}\cup\mathcal{U}$) ranges from 9% to 23%, indicating that a majority of predictions were Others. Among the LLMs, CodeLlama-7b achieved the highest AUP (23.4%), while StarCoder2-3b had the lowest overall AUP (9.3%). This may be attributed to the fact that StarCoder2 was not specifically fine-tuned on additional Python corpora, unlike other LLMs such as CodeLlama and CodeGen. Comparing the three versions of CodeGen suggests that the capacity of LLMs to predict plausible API usages increased with model size (i.e., 11.0%, 28.8%, and 23.6% for CodeGen-350m, -2b, and -6b, respectively), given the model architecture and training data remain consistent. However, it’s noteworthy that the largest LLM, GPT-3.5, did not achieve a high AUP (13.2%), possibly due to the instruction used not being finely tuned with advanced prompt engineering techniques such as chain-of-thought (COT) [65] and in-context learning (ICL) [66].
Summary 1: All the evaluated LLMs faced challenges in predicting plausible API usages, with AUP ranging from 10% to 30%. Effectiveness of LLM-based code completion generally improved with model size and language-specific fine-tuning.
Figure 3: Distribution of RepC, DepC, and Others Completions by LLMs for Prompts from $\mathcal{O}$ and $\mathcal{U}$
TABLE III: AUP and DUR for $\mathcal{O}$, $\mathcal{U}$, and $\mathcal{O}\cup\mathcal{U}$
@@
  CodeGen-350m 7.9 12.4 11.0 77.7 9.0 24.9
CodeGen-2b 15.9 20.1 18.8 89.6 11.3 32.6
CodeGen-6b 20.7 23.6 22.7 90.0 11.1 34.2
DeepSeek-1.3b 9.2 11.8 11.0 69.7 11.6 27.2
StarCoder-3b 8.1 9.9 9.3 79.4 17.0 34.4
CodeLlama-7b 21.6 24.3 23.4 86.9 12.1 34.2
GPT-3.5 11.9 13.8 13.2 85.7 17.8 37.4
@@
Summary 2: All the evaluated LLMs faced issues with deprecated API usages, with DUR ranging from 25% to 38%, and larger models exhibiting higher DUR. Among the LLMs, CodeLlama-7b and CodeGen-6b demonstrated the best balance between AUP and DUR.
IV-A2 Potential Cause Discussion
From the model perspective, the causes of deprecated API usages consist of two main points:
Model Training: As libraries evolve, both deprecated APIs and their replacements are commonly found in open-source code repositories. Since the training data for LLMs are primarily sourced from these repositories without filtering for deprecated APIs, they often include instances of deprecated API usage. While the training datasets for specific LLMs like CodeLlama and GPT-3.5 are not publicly available, preventing direct inspection, the distribution of deprecated API usages in open-source code repositories can serve as a proxy for understanding the nature of their training data. In this study, we identified 9,022 instances of deprecated API usage and 19,103 instances of their replacements in open-source code repositories. These coarse statistics suggest that a significant portion of the training data likely includes deprecated APIs. Moreover, prior studies have demonstrated a relationship between training data distribution and model output behavior [67, 68]. When trained on data containing deprecated APIs, LLMs may “memorize” these APIs and associated usage patterns as part of their learned knowledge [69, 70]. The different training datasets also led to different AUP and DUR of the LLMs.
Model Inference: LLMs generated completions by predicting token probabilities based on their learned prior knowledge (e.g., memorized API usage contexts) and applying token selection strategies (e.g., greedy search or beam search). Given certain contexts, LLMs were likely to predict deprecated API usages due to high token probabilities, without considering any posterior API deprecation knowledge. To support this hypothesis, we calculated the generative likelihoods of both $dep$ and $rep$ for each prompt $pmpt$ across open-source LLMs by analyzing predicted token probabilities during decoding. To begin, we input $pmpt$ into an LLM, followed by sequentially feeding each token in $dep$ (or $rep$) into the model. Throughout this process, we recorded the predicted token probability distributions prior to each token in $dep$ (or $rep$) being input. Next, we extracted the predicted probabilities for tokens in $dep$ (or $rep$) from these distributions using their vocabulary indices. The generative likelihoods of $dep$ and $rep$ were then computed by summing the log probabilities of their tokens, a common method in language modeling [71]. We performed a paired t-test on these likelihoods. The t-statistic values indicate that for all six open-source LLMs, the likelihoods of deprecated APIs are significantly higher than those of replacing APIs for prompts from $\mathcal{O}$, whereas the opposite is true for prompts from $\mathcal{U}$. The $p$ values are approximately zero.
Summary 3: There are two primary reasons why LLMs predict deprecated APIs: the presence of deprecated API usages in corpora during model training, and the absence of posterior knowledge about API deprecations during model inference.
IV-B RQ2: Prompt Perspective for Status Quo and Root Causes
@@ IV-B1 Status Quo Analysis
Table III presents the AUP and DUR for prompts from the two datasets, i.e., $\mathcal{O}$ and $\mathcal{U}$.
@@
API Usage Plausibility. Between the prompts from the two different datasets, $\mathcal{O}$ and $\mathcal{U}$, there are some differences in AUP for all LLMs, with relative differences (i.e., $(AUP^{\mathcal{U}}-AUP^{\mathcal{O}})/AUP^{\mathcal{O}}$) ranging from 12.5% to 60.0%. This disparity may be attributed to the imbalance in the number of outdated and up-to-dated functions in the LLMs’ training corpora [72, 73]. Indirect evidence for this is that, in this study, the up-to-dated functions collected from open-source code repositories were about twice as many as the outdated functions (i.e., 19,103 vs. 9,022), even though there was no collection preference. Given that LLMs were often trained on open-source code repositories, they likely learned more up-to-dated functions than outdated functions, leading to better AUP for the $\mathcal{U}$ dataset. Additionally, larger LLMs showcased smaller AUP differences (e.g., 16% for CodeLlama-7b), possibly due to the better generalizability.
@@
Summary 4: The LLMs showcased difference in AUP between the two datasets $\mathcal{O}$ and $\mathcal{U}$. This difference is possibly attributed to the different distribution of outdated and up-to-dated functions in the training corpora of LLMs.
@@
Deprecated Usage Rate. When considering $\mathcal{O}$ and $\mathcal{U}$ separately, all LLMs consistently demonstrated extremely high deprecated usage rates for $\mathcal{O}$ (70%-90% DUR) and relatively low rates for $\mathcal{U}$ (9%-18% DUR). This significant difference is also evident in the distribution of RepC and DepC completions shown in Figure 3. This observed discrepancy can be attributed to the tendency of LLMs to predict reference APIs (i.e., the APIs used in the original functions) for most prompts in both the $\mathcal{O}$ and $\mathcal{U}$ datasets, reflecting the unique contextual characteristics of each prompt. To quantify this, we define an auxiliary metric, the Reference Usage Rate (RUR), as follows:
$\text{RUR}=\frac{\sum_{p\in\mathcal{P}}\mathbb{I}(\textbf{LLM}(p)=reference)}{\sum_{p\in\mathcal{P}}\mathbb{I}(\text{anno}(\textbf{LLM}(p))\in\{\text{\emph{DepC}},\text{\emph{RepC}}\})},$
which measures the proportion of reference APIs in plausible completions. RUR is equal to DUR for $\mathcal{O}$ and $(1-DUR)$ for $\mathcal{U}$. Thus, no significant difference exists between $\mathcal{O}$ and $\mathcal{U}$ in terms of RUR, which is consistently within the 70%-90% range for both. Nonetheless, the 9%-18% DUR for $\mathcal{U}$ indicates that LLMs still predicted the usage of deprecated APIs, even for the prompts from up-to-dated functions.
@@
Summary 5: The LLMs consistently exhibited a significant difference in DUR between the two datasets, with extremely high deprecated API usage rates for $\mathcal{O}$ (70%-90% DUR) and relatively low rates for $\mathcal{U}$ (9%-18% DUR).
IV-B2 Potential Cause Discussion